OpenStudy (anonymous):

Solve the equation x3 = 8 A. 2 B. 2.7 C. 24 D. 512

OpenStudy (tkhunny):

Presume that is \(x^{3} = 8\) ?

OpenStudy (anonymous):

yes i couldnt get the format right

OpenStudy (anonymous):

can anybody help please

OpenStudy (tkhunny):

You can use the equation editor, below, to make expressions like that. Or, you can use the carat for exponentiation. x^3 = 8 means the same thing.

OpenStudy (tkhunny):

1*1*1 = 1 2*2*2 = 8 3*3*3 = 27 Didn't need the last one, but I threw it in for practice.

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

@kpl199907 could you ask for help somewhere else

OpenStudy (anonymous):

sorry yes

OpenStudy (anonymous):

lol i just tryna get my own help @kpl199907

OpenStudy (anonymous):

k that's cool

OpenStudy (studygurl14):

hey @kpl199907 I can help. open it in new thread

OpenStudy (jdoe0001):

\(\bf x^3=8\qquad {\color{blue}{ 8\implies 2\cdot 2\cdot 2\implies 2^3}}\\ \quad \\ x^3=2^3\qquad \large \textit{take }\sqrt[3]{\qquad }\textit{ to both sides, thus}\\ \quad \\ \Large \sqrt[3]{x^3}=\sqrt[3]{2^3}\)